# Mental Health Impacts of AI Companions

### v1

@@ 4.2. Matching and Causal Inference Framework
To examine changes in users’ psychosocial outcomes after interacting with AICCs, and mitigate confounds, we adopted a causal-inference design inspired by the potential outcomes framework (Rubin, 2005). Our goal is to examine the causal effects of engaging with AICCs (i.e., the treatment) compared to what would have happened if the treatment was not administered (i.e., the counterfactual). However, given that it is practically impossible to identify a true counterfactual, the potential outcomes framework guides the identification of comparable (or Control) individuals based on pre-treatment characteristics. This step is known as matching, and we adopted a stratified propensity score matching (Saha et al., 2019b; Yuan et al., 2023a; Lambert et al., 2024), which accounts for both online behavior and other observable traits (covariates), and enabled us to isolate potential treatment effects. If two individuals (generalizable to groups) were similar in relevant characteristics and external conditions, then any difference in outcome was plausibly associated with interaction with the AICC. Our approach of stratified propensity score matching allowed us to handle the bias-variance tradeoff by striking a balance between too biased (one-to-one matching) and too variant (unmatched) data comparisons so that we can isolate and examine the effects within each stratum (Saha et al., 2019b).
Covariates.  Our matching approach controlled for four categories of covariates. First, we incorporated Reddit behavioral features, including the number of posts, the number of comments, the number of days that users had at least one post or comment, and overall posting frequency, to account for general activity levels on Reddit. Second, we extracted the top 500 unigrams from users’ posts to capture salient lexical patterns indicative of individual language use. Third, we measured psycholinguistic traits using the LIWC lexicon, which reflected psychological states and linguistic markers related to affect, cognition, and interpersonal focus. Finally, we included the average frequency of posts expressing symptomatic mental health expressions (depression, anxiety, stress, suicidal ideation, loneliness), as well as measures of grief valence and grief activation. These covariates accounted for both behavioral and psychological baselines, enabling a more robust assessment of the psychosocial impact of AICC interactions.
@@ 4.2.1. Stratified Propensity Score Matching Approach
To ensure comparability across the three user groups (Treatment, Control-AI, and Control-NAI), we employed stratified propensity score matching. For this purpose, we implemented a binary logistic regression model to estimate the propensity of users belonging to the Treatment group versus either Control group. This model was trained using the four sets of covariates described earlier. The resulting propensity scores reflected the likelihood of being in the treatment group, given observed covariates. Propensity scores (011) were then divided into 100 equal-width strata, grouping users with similar scores within each stratum.. We excluded users with propensity scores outside two standard deviations from the mean and dropped strata that did not contain at least 5 users in both the treatment and corresponding control group (figure LABEL:sfig:a_matching). After filtering, we retained 70 strata, comprising 1,984 Treatment users, 5,657 Control-AI users, and 6,770 Control-NAI users.
@@ 4.3. Difference-in-Differences Regression Analysis
To further enhance the robustness of our estimation of the causal effect of AICC engagement on psychosocial outcomes, we adopted a DiD regression approach (Lambert et al., 2025). Our matching procedure controlled for observable covariates (e.g, posting activity). The DiD method further addressed time-invariant unobserved confounding, though it could not fully account for time-varying external shocks (e.g., global news) that may have influenced psychosocial expression (Lambert et al., 2025). We applied DiD analysis separately for the following comparisons: (1) Treatment vs. Control-AI users, and (2) Treatment vs. Control-NAI users. This allowed us to assess whether observed effects were specific to AICC interaction or generalizable to broader platform engagement. For each user, we calculated the average value of each outcome (excluding the active days) in the pre- and post-treatment/placebo windows—one year before and one year after the treatment event. To support valid causal inference, DiD analysis required the parallel trends assumption (Lechner et al., 2011). We empirically verified this assumption, with results presented in the Appendix.
@@ 4.4. Findings: Psychosocial Changes After AICC Engagement
Table 4. Summary of psychosocial outcome differences between Treatment (Tr.) and two Control groups: Control-AI (Ct-AI) and Control-NAI (Ct-NAI). For each outcome, we report the mean difference ($\Delta$), Average Treatment Effect (ATE), ATE percentage, Cohen’s $d$, $t$-statistic, and the coefficient of the $a_{t}*\delta(d\geq 0)$ interaction term (DiD) in DiD regression results. All $p$-values are Bonferroni-corrected for multiple comparisons. The DiD column presents results from the DiD regression analysis, capturing post-treatment relative changes in psychosocial outcomes. Statistical significance reported after Bonferroni corrections: * pp¡0.05, ** pp¡0.01, ***pp¡0.001. Statistically significant values are bold-faced.:
Treatment Control-AI Control-NAI
$\Delta$ $\Delta$ ATE (%) $\mathtt{d}$ $\mathtt{t}$ DiD (%) $\Delta$ ATE (%) $\mathtt{d}$ $\mathtt{t}$ DiD (%)
@@
LIWC: Affective 0.00 -0.10 13.00 0.12 3.30 * 0.23 *** -0.05 5.46 0.02 0.69 0.30
Grief: Valence 0.20 0.04 15.77 0.42 13.07 *** 7.58 *** 0.16 6.52 0.22 6.47 *** -2.54
Grief: Activation 0.21 0.04 17.72 0.45 13.98 *** 6.02 *** 0.16 7.67 0.23 7.01 *** -2.91
Loneliness 0.03 -0.47 105.92 0.19 4.19 ** 0.58 *** -0.30 63.26 0.14 3.60 ** 0.44 ***
DASS: Depression 0.08 0.18 -4.80 -0.04 -1.32 1.83 *** -0.04 14.99 0.18 5.67 *** 1.67 ***
DASS: Anxiety 0.02 0.33 -20.27 -0.18 -6.14 *** 0.28 -0.06 9.60 0.09 2.87 0.54 ***
DASS: Stress 0.08 0.11 -1.27 0.03 0.9 1.35 *** 0.02 9.60 0.16 5.16 *** 2.12 ***
DASS: Suicidal ideation 0.12 -0.10 28.28 0.12 3.72 ** 0.95 *** -0.17 37.67 0.19 5.82 *** 0.61 ***
@@
Interactivity 0.07 0.33 -18.29 -0.13 -5.29 *** -0.58 ** -0.05 8.71 0.12 4.89 *** 0.09
Topical Diversity 0.00 0.00 -0.44 -0.10 -3.36 * -0.21 *** 0.00 -0.18 -0.05 -1.67 0.19 ***
Posting Frequency 0.79 0.25 33.00 0.24 6.84 *** 49.66 *** 0.83 -3.01 0.00 0.00 -12.79 *
Active Days 0.66 0.31 22.09 0.24 8.42 *** NA 0.76 -2.91 -0.01 -0.35 NA
@@
Readability 0.08 0.02 4.59 0.20 5.82 *** 149.28 *** 0.06 2.68 0.06 2.74 -20.04
Complexity 0.00 0.02 -1.89 -0.06 -2.05 -14.97 ** 0.00 0.20 -0.01 -0.32 -11.71
Repeatability -0.06 -0.08 6.26 -0.02 -0.67 0.56 0.29 0.55 0.03 1.20 0.21
LIWC: Cog. & Percep. 0.07 0.00 8.03 0.22 7.21 *** -0.08 -0.02 8.74 0.23 7.79 *** 0.32 ***
LIWC: Social Context 0.00 -0.04 3.67 0.08 2.58 -0.28 * -0.03 1.27 0.00 0.14 -0.18
LIWC: Lexical Density 0.03 0.00 2.93 0.19 6.56 *** -0.55 *** 0.00 3.19 0.14 5.51 *** 0.42 *
LIWC: Interpersonal Focus 0.10 -0.03 13.47 0.29 9.15 *** 0.74 *** -0.04 12.89 0.21 8.10 *** 0.76 ***
LIWC: Temporal References 0.08 -0.02 10.20 0.32 10.02 *** 0.07 -0.02 8.94 0.24 8.05 *** 0.24 ***
@@ 4.4.1. Affective Outcomes
As shown in Table 4, we find significant affective changes among Treatment users after engaging with AICCs. The Treatment users exhibited a statistically significant increase when compared to Control-AI users (ATE = 13.00%), suggesting heightened emotional expressiveness following engagement. In contrast, the difference between the Treatment group and Control-NAI was not statistically significant. Results from the DiD regression further corroborated these findings: the post-treatment change in affective word usage was modest but significant relative to Control-AI (DiD = 0.23%), whereas the increase relative to Control-NAI (DiD = 0.30%) did not show statistical significance.
@@
More pronounced effects were observed in grief-related dimensions. Specifically, for grief-valence, Treatment users exhibited substantial increases compared to both Control groups (Control-AI: ATE=15.77%; Control-NAI: ATE = 6.52%). DiD regression further confirmed a significant post-treatment increase in grief-valence relative to Control-AI (DiD=7.58%), but not to Control-NAI. Similarly, Treatment users showed increased grief-activation compared to Control-AI and Control-NAI users (Control-AI: ATE = 17.71%; Control-NAI: ATE=7.67%), with DiD indicating a positive effect compared to Control-AI (6.02%).
@@
Symptomatic expressions of depression, anxiety, and stress revealed mixed treatment effects with contrasting patterns across Control groups. The use of symptomatic expressions of depression by Treatment users showed non-significant decreases relative to Control-AI users but substantial increases compared to Control-NAI users (ATE=14.99%). For anxiety expressions, Treatment users showed a significant decrease compared to Control-AI users (ATE=-20.27%) but non-significant effects relative to Control-NAI. The use of stress expressions by Treatment users demonstrated non-significant results against Control-AI users, but moderate positive effects compared to Control-NAI (ATE=9.60%). DiD regression consistently reveals significant post-treatment increases across stress and depression measurements relative to both control groups (depression:Control-NAI DiD=1.67%; stress: Control-AI DiD=1.35%, Control-NAI DiD =2.12%).
@@
Symptomatic expressions of suicide ideation and loneliness showed the most pronounced and concerning effects. The symptomatic expressions of suicide ideation in the Treatment group showed substantial increases compared to both control groups (Control-AI: ATE=28.28%; Control-NAI: ATE=37.67%). DiD regression confirms significant increases relative to both control groups (Control-AI: DiD = 0.95%; Control-NAI: DiD = 0.61%). Similarly, symptomatic expressions of loneliness showed substantial negative effects, with treatment group participants demonstrating marked increases compared to Control-AI (ATE = 105.92%) and Control-NAI (ATE = 63.26%), with DiD confirming significant post-treatment rises across groups.
@@ 4.4.2. Behavioral Outcomes
Behavioral outcomes revealed complex patterns of interaction and engagement. Treatment group users demonstrated decreased interactivity compared to Control-AI users (ATE = -18.29%; DiD=-0.58%), possibly suggesting more introspective engagement. Conversely, interactivity increased slightly compared to Control-NAI users (ATE=8.71%), but no significant post-treatment DiD difference was observed.
Topical diversity decreased significantly among Treatment users compared to Control-AI users (ATE=-0.44%; DiD = -0.21%), with no significant change relative to Control-NAI users. Posting frequency increased significantly relative to Control-AI users (ATE=33.00%; DiD=49.66%), suggesting enhanced engagement on Reddit, but no significant change relative to Control-NAI users.
@@ 4.4.3. Cognitive Outcomes
Treatment individuals showed statistically improved readability compared to Control-AI users (ATE=4.59% DiD=149.28%), reflecting increased clarity and better structured language in expression, plausibly indicating cognitive enhancement or emotional recovery (Saha et al., 2018; Ernala et al., 2017).
Complexity among Treatment users slightly decreased relative to Control-AI users; however, statistical tests showed no significant differences compared with either Control-AI or Control-NAI users. Repeatability marginally increased relative to Control-AI and Control-AI users, but statistical tests revealed no significant changes. For repeatability among Treatment users, no significant differences were observed relative to Control-AI and Control-NAI users, suggesting limited impact on cognitive flexibility or expressive variability.
Within LIWC categories, Treatment users showed an increase in cognition and perception compared to Control-NAI (ATE=8.74%, dd=0.23), with a significant difference also found relative to Control-AI. However, DiD regression showed no significant difference. Likewise, Treatment users also showed an increase in lexical density and awareness relative to both Control-AI (dd=0.19) and Control-NAI (dd=0.14). Treatment users also showed an increase in interpersonal focus with strong significance across both comparisons (Control-AI and Control-NAI). Treatment users’ temporal references also increased significantly relative to Control-NAI (dd=0.24), suggesting improved temporal awareness in communication.

### v2

@@ 4.2. Matching and Causal Inference Framework
Ideally, the effects of an intervention—in our case, the use of AICCs—would be best assessed through a randomized controlled trial (RCT), where users are randomly assigned to engage (or not engage) with an AICC. However, in the context of sensitive, voluntary, and personalized technologies like AICCs, randomization is neither practical nor ethical. Users choose to adopt these tools based on their own needs and circumstances, and any attempt to impose artificial exposure would undermine ecological validity and user autonomy, and introduce biases such as the observer effect (Saha et al., 2024). In contrast, our quasi-experimental design leverages naturalistic and observational data from users who self-select into AICC use, enabling the examination of real-world psychosocial trajectories without disrupting authentic engagement.
In our quasi-experimental design, to examine changes in users’ psychosocial outcomes after interacting with AICCs, and mitigate confounds, we adopted a causal-inference design inspired by the potential outcomes framework (Rubin, 2005; Imbens and Rubin, 2015). Our goal is to examine the causal effects of engaging with AICCs (i.e., the treatment) compared to what would have happened if the treatment was not administered (i.e., the counterfactual).
Given that it is practically impossible to identify a true counterfactual, the potential outcomes framework guides the identification of comparable (or Control) individuals based on pre-treatment characteristics. This step is known as matching, and we adopted a stratified propensity score matching (Saha et al., 2019c; Yuan et al., 2023a; Lambert et al., 2024), which accounts for online behavior and observable traits (covariates), and enabled us to isolate potential treatment effects. If two individuals (generalizable to groups) were similar in relevant characteristics and external conditions, then any difference in outcome was plausibly associated with interaction with the AICC. Our approach of stratified propensity score matching allowed us to handle the bias-variance tradeoff by striking a balance between too biased (one-to-one matching) and too variant (unmatched) data comparisons so that we can isolate and examine the effects within each stratum (Saha et al., 2019c).
Covariates.  Our matching approach controlled for four covariate categories. First, we incorporated Reddit behavioral features, including the number of posts, the number of comments, the number of days that users had at least one post or comment, and overall posting frequency, to account for general activity levels on Reddit. Second, we extracted the top 500 unigrams from users’ posts to capture salient lexical patterns indicative of individual language use. Third, we measured psycholinguistic traits using the LIWC lexicon, which reflected psychological states and linguistic markers related to affect, cognition, and interpersonal focus. Finally, we included the average frequency of posts expressing symptomatic mental health expressions (depression, anxiety, stress, suicidal ideation, loneliness), as well as measures of grief valence and grief activation. These covariates accounted for both behavioral and psychological baselines, enabling a more robust assessment of the psychosocial impact of AICC interactions.
@@ 4.2.1. Stratified Propensity Score Matching Approach
To ensure comparability across the three user groups (Treatment, Control-VA, and Control-NAI), we employed stratified propensity score matching. For this purpose, we implemented a binary logistic regression model to estimate the propensity of users belonging to the Treatment group versus either Control group. This model was trained using the four sets of covariates described earlier. The resulting propensity scores reflected the likelihood of being in the treatment group, given observed covariates. Propensity scores (011) were then divided into 100 equal-width strata, grouping users with similar scores within each stratum.. We excluded users with propensity scores outside two standard deviations from the mean and dropped strata that did not contain at least 5 users in both the treatment and corresponding control group (figure LABEL:sfig:a_matching). After filtering, we retained 70 strata, comprising 1,984 Treatment users, 5,849 Control-LLM users, 5,657 Control-VA users, and 6,770 Control-NAI users.
@@ 4.3. Difference-in-Differences Regression Analysis
To further enhance the robustness of our estimation of the causal effect of AICC engagement on psychosocial outcomes, we adopted a DiD regression approach (Lambert et al., 2025). Our matching procedure controlled for observable covariates (e.g, posting activity). The DiD method further addressed time-invariant unobserved confounding, though it could not fully account for time-varying external shocks (e.g., global news) that may have influenced psychosocial expression (Lambert et al., 2025). We applied DiD analysis separately for the following comparisons: (1)Treatment vs. Control-LLM users, (2) Treatment vs. Control-VA users, and (3) Treatment vs. Control-NAI users. This allowed us to assess whether observed effects were specific to AICC interaction or generalizable to broader platform engagement. For each user, we calculated the average value of each outcome (excluding the active days) in the pre- and post-treatment/placebo windows—one year before and one year after the treatment event. To support valid causal inference, DiD analysis required the parallel trends assumption (Lechner et al., 2011). We empirically verified this assumption, with results presented in the Appendix.
@@ 4.4. Findings: Psychosocial Changes After AICC Engagement
Table summarizing psychosocial outcome differences between treatment and three control groups. Outcomes are grouped into affective, behavioral, and cognitive categories, such as affective language, grief, loneliness, depression, anxiety, suicidal ideation, interactivity, topical diversity, posting frequency, readability, complexity, and linguistic markers. For each outcome, the table reports mean differences, average treatment effects, test statistics, and DiD regression coefficients. Statistically significant results after Bonferroni correction are bolded, highlighting notable changes across affective, behavioral, and cognitive outcomes.
Table 4. Summary of psychosocial outcome differences between Treatment (Tr.) and three Control groups: Control-LLM, Control-VA and Control-NAI. For each outcome, we report the mean difference ($\Delta$), Average Treatment Effect (ATE) percentage, $t$-statistic, and the coefficient of the $a_{t}*\delta(d\geq 0)$ interaction term (DiD) in DiD regression results. All $p$-values are Bonferroni-corrected for multiple comparisons. The DiD column presents results from the DiD regression analysis, capturing post-treatment relative changes in psychosocial outcomes. Statistical significance reported after Bonferroni corrections: * pp¡0.05, ** pp¡0.01, ***pp¡0.001. Statistically significant values are bold-faced.
Tr. Control-LLM Control-VA Control-NAI
$\Delta$ $\Delta$ ATE (%) $\mathtt{t}$ DiD (%) $\Delta$ ATE (%) $\mathtt{t}$ DiD (%) $\Delta$ ATE (%) $\mathtt{t}$ DiD (%)
@@
LIWC: Affective 0.00 -0.12 11.24 2.00 ** 0.27 *** -0.10 13.00 3.30 * 0.23 *** -0.05 5.46 0.69 0.30
Grief: Valence 0.20 0.07 2.17 5.07 *** 6.68 *** 0.04 15.77 13.07 *** 7.58 *** 0.16 6.52 6.47 *** -2.54
Grief: Activation 0.21 0.07 3.45 5.50 *** 4.7 ** 0.04 17.72 13.98 *** 6.02 *** 0.16 7.67 7.01 *** -2.91
Loneliness 0.03 -0.43 91.02 3.64 ** 0.58 *** -0.47 105.92 4.19 ** 0.58 *** -0.30 63.26 3.60 ** 0.44 ***
DASS: Depression 0.08 0.03 6.36 3.05 * 1.28 *** 0.18 -4.80 -1.32 1.83 *** -0.04 14.99 5.67 *** 1.67 ***
DASS: Anxiety 0.02 0 2.42 1.31 0.68 *** 0.33 -20.27 -6.14 *** 0.28 -0.06 9.60 2.87 0.54 ***
DASS: Stress 0.08 0.06 2.1 2.93 0.65 * 0.11 -1.27 0.9 1.35 *** 0.02 9.60 5.16 *** 2.12 ***
DASS: Suicidal ideation 0.12 -0.19 51.89 5.19 *** 1.39 *** -0.10 28.28 3.72 ** 0.95 *** -0.17 37.67 5.82 *** 0.61 ***
@@
Interactivity 0.07 0.12 -10.62 -0.99 -0.54 ** 0.33 -18.29 -5.29 *** -0.58 ** -0.05 8.71 4.89 *** 0.09
Topical Diversity 0.00 0 -0.36 -1.54 -0.23 *** 0.00 -0.44 -3.36 * -0.21 *** 0.00 -0.18 -1.67 0.19 ***
Posting Frequency 0.79 0.62 17.7 3.56 ** -30.23 *** 0.25 33.00 6.84 *** 49.66 *** 0.83 -3.01 0.00 -12.79 *
Active Days 0.66 0.44 24.18 7.29 *** NA 0.31 22.09 8.42 *** NA 0.76 -2.91 -0.35 NA
@@
Readability 0.08 0.12 -5.52 -1.34 -8.05 0.02 4.59 5.82 *** 149.28 *** 0.06 2.68 2.74 -20.04
Complexity 0.00 -0.1 -8.45 -3.05 * -192.77 *** 0.02 -1.89 -2.05 -14.97 ** 0.00 0.20 -0.32 -11.71
Repeatability -0.06 0.24 3.02 1.44 -0.41 -0.08 6.26 -0.67 0.56 0.29 0.55 1.20 0.21
LIWC: Cog. & Percep. 0.07 0.02 4.89 3.21 * -0.16 0.00 8.03 7.21 *** -0.08 -0.02 8.74 7.79 *** 0.32 ***
LIWC: Social Context 0.00 0.01 0.78 -1.23 -0.23 * -0.04 3.67 2.58 -0.28 * -0.03 1.27 0.14 -0.18
LIWC: Lexical Density 0.03 0 3.88 4.73 *** -0.12 0.00 2.93 6.56 *** -0.55 *** 0.00 3.19 5.51 *** 0.42 *
LIWC: Interpersonal Focus 0.10 0 10.22 6.09 *** 0.48 *** -0.03 13.47 9.15 *** 0.74 *** -0.04 12.89 8.10 *** 0.76 ***
LIWC: Temporal References 0.08 -0.01 9.94 6.25 *** 0.07 -0.02 10.20 10.02 *** 0.07 -0.02 8.94 8.05 *** 0.24 ***
@@ 4.4.1. Affective Outcomes
As shown in Table 4, our results revealed significant affective changes among Treatment users after engaging with AICCs.The usage of affective words among Treatment users slightly increased relative to Control-LLM users The Treatment users exhibited a statistically significant increase when compared to Control-VA users (ATE = 13.00%), suggesting heightened emotional expressiveness following engagement. In contrast, the difference between the Treatment group and Control-NAI was not statistically significant. Results from the DiD regression confirmed the post-treatment change in affective word usage was modest but significant relative to Control-VA (DiD = 0.23%), whereas the increase relative to Control-LLM and Control-NAI did not show statistical significance.
@@
More pronounced effects were observed in grief-related dimensions. Specifically, for grief-valence, Treatment users exhibited substantial increases compared to all Control groups (Control-LLM: ATE=2.17%, Control-VA: ATE=15.77%; Control-NAI: ATE = 6.52%). DiD regression further confirmed a significant post-treatment increase in grief-valence relative to Control-LLM (DiD=6.68%) and Control-VA (DiD=7.58%), but not to Control-NAI. Similarly, Treatment users showed increased grief-activation compared to Control-LLM, Control-VA and Control-NAI users (Control-LLM: ATE=3.45, Control-VA: ATE = 17.71%; Control-NAI: ATE=7.67%), with DiD indicating a positive effect compared to Control-LLM (4.70) and Control-VA (6.02%).
@@
Symptomatic expressions of anxiety, stress, and depression revealed mixed treatment effects with contrasting patterns across Control groups. The use of symptomatic expressions of depression by Treatment users showed non-significant decreases relative to Control-VA users but substantial increases compared to Control-LLM (ATE=3.05%) and Control-NAI users (ATE=14.99%). For symptomatic expressions of anxiety, Treatment users showed a significant decrease compared to Control-VA users (ATE=-20.27%) but non-significant effects relative to Control-LLM and Control-NAI. The use of symptomatic expressions of stress by Treatment users demonstrated non-significant results against Control-LLM and Control-VA users, but moderate positive effects compared to Control-NAI (ATE=9.60%). DiD regression consistently reveals significant post-treatment increases across stress and depression measurements relative to all control groups.
@@
Symptomatic expressions of suicide ideation and loneliness showed the most pronounced and concerning effects. The symptomatic expressions of suicide ideation in the Treatment group showed substantial increases compared to all control groups (Control-LLM: ATE=51.89%, Control-VA: ATE=28.28%; Control-NAI: ATE=37.67%). DiD regression confirms significant increases relative to both control groups (Control-LLM: DID=1.39%, Control-VA: DiD = 0.95%; Control-NAI: DiD = 0.61%). Similarly, symptomatic expressions of loneliness showed substantial negative effects, with treatment group participants demonstrating marked increases compared to Control-LLM: ATE=91.02%, Control-VA (ATE = 105.92%) and Control-NAI (ATE = 63.26%), with DiD confirming significant post-treatment rises across groups.
@@ 4.4.2. Behavioral Outcomes
Behavioral outcomes revealed complex patterns of interaction and engagement. Treatment group users demonstrated decreased interactivity compared to Control-VA users (ATE=-18.29%; DiD=-0.58%), possibly suggesting more introspective engagement. Conversely, interactivity increased slightly compared to Control-NAI users (ATE=8.71%), but no significant post-treatment DiD difference was observed.
Topical diversity decreased significantly among Treatment users compared to Control-LLM (ATE=-0.99%; DiD=-0.23%) and Control-VA users (ATE=-0.44%; DiD = -0.21%), with no significant change relative to Control-NAI users. Posting frequency increased significantly relative to Control-VA users (ATE=33.00%; DiD=49.66%), suggesting enhanced engagement on Reddit, but no significant change relative to Control-NAI users.
@@ 4.4.3. Cognitive Outcomes
Treatment individuals showed statistically improved readability compared to Control-VA users (ATE=4.59% DiD=149.28%), reflecting increased clarity and better structured language in expression, plausibly indicating cognitive enhancement or emotional recovery (Saha et al., 2018; Ernala et al., 2017).
Complexity among Treatment users slightly decreased relative to Control-LLM users; however, statistical tests showed no significant differences compared with either Control-VA or Control-NAI users. Repeatability marginally increased relative to Control-LLM, Control-VA and Control-VA users, but statistical tests revealed no significant changes, suggesting limited impact on cognitive flexibility or expressive variability.
Within LIWC categories, Treatment users showed an increase in cognition and perception compared to Control-NAI (ATE=8.74%, dd=0.23), with a significant difference also found relative to Control-LLM and Control-VA. However, DiD regression showed no significant difference. Likewise, Treatment users also showed a significant increase in lexical density and awareness relative to both Control-VA and Control-NAI. Treatment users also showed an increase in interpersonal focus with strong significance across all comparisons (Control-LLM, Control-VA and Control-NAI). Treatment users’ temporal references also increased significantly relative to Control-NAI, suggesting improved temporal awareness in communication.